ID |
Date |
Author |
Type |
Category |
Subject |
10197
|
Mon Jul 14 17:51:34 2014 |
Nichin | Update | Computer Scripts / Programs | MEDM for PDFR system |
A
Successfully completed the rudimentary GUI for PDFR system. (users/nichin/PDFR)
Pressing any of the buttons above runs the script that does the following:
1) Change RF mux channel to the required one.
2) Frequency sweep on the network analyzer. The common sweep parameters are in a file named param_NWAG4395A.yml . PD specific parameters are in param_[PD name].yml in their respective folders
3) The transimpedance is calculated and the plot is saved as PDF in the respective folder for the PD. Each set of measurement data and plot is in a timestamped subfolder.
Further work:
To take transimpedance readings for each PD and create a canonical set of data that can be used to compare with data obtained for every measurement run. |
10196
|
Mon Jul 14 16:51:07 2014 |
Nichin | Update | Electronics | Martian table updated, Named server restarted |
[Nichin, Jenne]
The martian lookup tables are located at /etc/bind/zones/martian.db and etc/bind/zones/rev.113.168.192.in-addr.arpa
Jenne updated these to include santuzza.martian 192.168.113.109
The method to restart named server given at https://wiki-40m.ligo.caltech.edu/Martian_Host_Table also does not work.
I restarted it using >sudo /etc/init.d/bind9 restart
The named server is now updated and works fine. :) I will update the 40m wiki now. |
10195
|
Mon Jul 14 16:19:41 2014 |
Andres | Update | 40m Xend Table upgrade | Took the measurement for the Mode Matching |
Nick and I measured the reflected power of the green light in locked and unlocked. I'm working on the calculation of the mode matching. Tonight, I'll be posted my calculation I'm still working on it.
JCD: Andres forgot to mention that they closed the PSL shutter, so that they could look at the green light that is reflected off the harmonic separator toward the IR trans path. Also, the Xarm (and the Yarm) were aligned to IR using the ASS, and then ASX was used to align the green beam to the cavity. |
10194
|
Mon Jul 14 14:28:27 2014 |
Koji | Summary | Electronics | Timing Issues of Mini Circuits UFC-6000: Solved |
Looks good. Now you have the internal timer to verify the external clock.
If you can realize the constant rate sampling without employing the external clock, that's quite handy. |
10193
|
Mon Jul 14 13:03:23 2014 |
Akhil | Summary | Electronics | Timing Issues of Mini Circuits UFC-6000: Solved |
Main Problem:
The frequency counter (FC) takes in an analog RF input(signal) and outputs the frequency of the signal(Ranging from 1 MHz- 6000 MHz) in the digital domain (into a processor). The FC samples the data with a given sample rate( user defined) which ranges from 0.1 s to 1 s(faced problems in fixing this initially). For data acquisition, we have been using a Raspberry Pi(as a processor) which is connected to the martian network and can communicate with the computers inside the 40m. The ultimate challenge which I faced( and been knocking my head off from past two-three weeks) is the synchronization of clocks between the Raspberry Pi and the FC i.e the clock which the FC uses to sample and dump data( every 'x' s) and the clock inside the raspberry pi( used in the loop to wait for a particular amount of time the frequency counter takes to dump successive data).
Steps Taken:
- To address this problem, first I added an external clock circuit which monitors the Raspberry Pi and the FC to dump and read data at a particular rate(which is equal to the sampling rate of the FC)In detail: http://nodus.ligo.caltech.edu:8080/40m/10129.
- While doing so, at first the level trigger algorithm was used which means that the external clock frequency was half as that of the reciprocal of the sampling rate and a trigger was seen every time the level shifts from +DC to -DC(of the external square wave).
- But this did not completely mitigate the issues and there were still few issues on how quickly the ADC reads the signal and R Pi processes it.
- To minimize these issues completely, an edge trigger algorithm which detects a pos edge(rising) of the clock was used. The clock frequency is now equal to the reciprocal of the sampling rate. This algorithm showed better results and greatly minimized the drift of the sampling time.
Psuedo Code(code attached):
open device : FC via USB-HID;
open device : ADC via I2C;
always(for t= recording time):
read data from ADC(external clock);
if pos edge detected:
read data from FC and store it in a register;
else read data from ADC;
end
write data stored in the register to a file( can be an Epics channel or a text file);
Results:
The attached are the plots showing the time between samples for a large number of samples taken for different sampling times of the FC. The percentage error is the percentage of standard error in the timing between two samples for the data for the entire measurement. It can be inferred that this error has been cut down to the order of ms.
To do next:
- I have started taking phase measurements( analysis and plots will follow this elog) and also the PSD plots with the improved timing characteristics.
|
Attachment 1: 0.2timinganalysis.png
|
|
Attachment 2: 0.3timinganalysis.png
|
|
Attachment 3: 0.5timinganalysis.png
|
|
Attachment 4: 1stiminganalysis.png
|
|
Attachment 5: pdf.zip
|
10192
|
Mon Jul 14 12:49:07 2014 |
Nichin | Update | Electronics | New Prologix GPIB-Ethernet controller |
Quote: |
Quote: |
I have configured a NEW Prologix GPIB-Ethernet controller to use with HP8591E Spectrum analyzer that sits right next to the control room computers.
Static IP: 192.168.113.109
Mask: 255.255.255.0
Gateway: 192.168.113.2
I have no clue how to give it a name like "something.martian" and to update the martian host table (Somebody please help!!)
|
The instructions for adding a name to the martian DNS table are in the wiki page that I pointed you to:
https://wiki-40m.ligo.caltech.edu/Martian_Host_Table
|
The instructions at https://wiki-40m.ligo.caltech.edu/Martian_Host_Table are outdated!
The name server configuration is currently at /etc/bind/zones/martian.db [ source: elog:10067 ]
Anyway, I need superuser access to edit the files, which I don't have. Even if I did know the password, I don't think it's a good idea for me to be messing around. So any of the 40m folks please update the martian table to include:
santuzza.martian 192.168.113.109
|
10191
|
Sun Jul 13 17:06:35 2014 |
Andres | Update | 40m Xend Table upgrade | Xarm Table Upgrade Calculation and Diagrams of possible new table layout |
Current Mode Matching and Gouy Phase Between Steering Mirrors
We found in 40m elog ID 3330 ( http://nodus.ligo.caltech.edu:8080/40m/3330) a documentation done by Kiwamu, where he measured the waist of the green. The waist of the green is about 35µm. Using a la mode, I was able to calculate the current mode matching, and the Gouy phase between the steering mirrors. In a la mode, I used the optical distances,which is just the distance measured times its index of refraction. I contacted someone from ThorLabs (which is the company that bought Optics For Research), and that person told that the Faraday IO-5-532-LP has a Terbium Gallium Garnet crystal of a length of 7mm and its index of refraction is 1.95. The current mode matching is 0.9343, and the current Gouy phase between steering mirrors is 0.2023 degrees. On Monday, Nick and I are planning to measure the actual mode matching. The attached below is the current X-arm optical layout.
Calculation For the New Optical Layout
Since the current Gouy phase between the steering mirror is essentially zero, we need to find a way how to increase the Gouy Phase. We tried to add two more lenses after the second steering mirror, and we found that increasing the Gouy phase result in a dramatically decrease in mode matching. For instance, a Gouy phase of about 50 degrees results in a mode matching of about .2, which is awful. We removed the first lens after the faraday, and we added two more mirrors and two more lenses after the second steering mirror. I modified the photo that I took and I place where the new lenses and new mirrors should go as shown in the second pictures attached below. Using a la mode, we found the following solution:
label z (m) type parameters
----- ----- ---- ----------
lens 1 0.0800 lens focalLength: 0.1000
First mirror 0.1550 flat mirror none:
Second mirror 0.2800 flat mirror none:
lens 2 0.4275 lens focalLength: Inf
lens 3 0.6549 lens focalLength: 0.3000
lens 4 0.8968 lens focalLength: -0.250
Third mirror 1.0675 flat mirror none:
Fourth mirror 1.4183 flat mirror none:
lens 5 1.6384 lens focalLength: -0.100
Fifth mirror 1.7351 flat mirror none:
Sixth mirror 2.0859 flat mirror none:
lens 6 2.1621 lens focalLength: 0.6000
ETM 2.7407 lens focalLength: -129.7
ITM 40.5307 flat mirror none:
The mode matching is 0.9786. The different Gouy phase different between Third Mirror and Fourth Mirror is 69.59 degrees, Gouy Phase between Fourth and Fifth 18.80 degrees, Gouy phase between Fifth and Sixth mirrors is 1.28 degrees, Gouy phase between Third and Fifth 88.38 degrees, and the Gouy phase between Fourth and Sixth is 20.08 degrees. Bellow attached the a la Mode code and the Plots.
Plan for this week
I don't think we have the lenses that we need for this new setup. Mostly, we will need to order the lenses on Monday. As I mention, Nick and I are going to measure the actual mode matching on Monday. If everything look good, then we will move on and do the Upgrade.
|
Attachment 1: CurrentOpticalLayout.png
|
|
Attachment 2: NewSetUp.PNG
|
|
Attachment 3: AlaModeSolutionplots.png
|
|
Attachment 4: EntireScaleRangeAlaModeSolution.png
|
|
Attachment 5: NewXarmOptimizationFromFaraday.m
|
close all
clear all
% In this code we are using a la mode to optimatize the mode matching and
% to optimatize the Gouy phase between mirror 1 and mirror 2. All the units
% are in meter
w0=(50*1e-6)/sqrt(2); % The Waist of the laser measured after SHG
z0_laser=-0.0083; % position measured where the waist is located
lamb= 532*10^-9; % wavelength of green light in mm
lFaraday=.0638; % Length of the faraday
... 209 more lines ...
|
10190
|
Sun Jul 13 11:37:36 2014 |
Jamie | Update | Electronics | New Prologix GPIB-Ethernet controller |
Quote: |
I have configured a NEW Prologix GPIB-Ethernet controller to use with HP8591E Spectrum analyzer that sits right next to the control room computers.
Static IP: 192.168.113.109
Mask: 255.255.255.0
Gateway: 192.168.113.2
I have no clue how to give it a name like "something.martian" and to update the martian host table (Somebody please help!!)
|
The instructions for adding a name to the martian DNS table are in the wiki page that I pointed you to:
https://wiki-40m.ligo.caltech.edu/Martian_Host_Table |
10189
|
Fri Jul 11 22:28:34 2014 |
Chris | Update | CDS | c1auxex "Unknown Host" |
Quote: |
c1auxex has forgotten who it is. Slow sliders for the QPD head were not responding, so I did a soft reboot from telnet. The machine didn't come back, so I plugged the RJ45-DB9 cable into the machine and looked at it through a minicom session. When I key the crate, it gives me an error that it can't load a file, with the error code 0x320001. Looking that up on a List of VxWorks error codes, I see that it is: S_hostLib_UNKNOWN_HOST (3276801 or 0x320001)
I'm not sure how this happened. I unplugged and replugged in the ethernet cable on the computer, but that didn't help. Rana is going in to wiggle the other end of the ethernet cable, in case that's the problem. EDIT: Replacing the ethernet cable did not help.
Former elogs that are useful: 10025, 10015
EDIT: The actual error message is:
boot device : ei
processor number : 0
host name : chiara
file name : /cvs/cds/vw/mv162-262-16M/vxWorks
inet on ethernet (e) : 192.168.113.59:ffffff00
host inet (h) : 192.168.113.104
user (u) : controls
flags (f) : 0x0
target name (tn) : c1auxex
startup script (s) : /cvs/cds/caltech/target/c1auxex/startup.cmd
Attaching network interface ei0... done.
Attaching network interface lo0... done.
Loading...
Error loading file: errno = 0x320001.
Can't load boot file!!
|
We fixed this problem (at least for now) by adding c1auxex to the /etc/hosts file on chiara (following a hint from this page). The DNS setup might be the culprit here. |
10188
|
Fri Jul 11 22:02:52 2014 |
Jamie, Chris | Omnistructure | CDS | cdsutils: multifarious upgrades |
To make the latest cdsutils available in the control room, we've done the following:
Upgrade pianosa to Ubuntu 12 (cdsutils depends on python2.7, not found in the previous release)
- Enable distribution upgrades in the Ubuntu Software Center prefs
- Check for updates in the Update Manager and click the big "Upgrade" button
Note that rossa remains on Ubuntu 10 for now.
Upgrade cdsutils to r260
- Instructions here
- cdsutils-238 was left as the default pointed to by the cdsutils symlink, for rossa's sake
Built and installed the nds2-client (a cdsutils dependency)
- Checked out the source tree from svn into /ligo/svncommon/nds2
- Built tags/nds2_client_0_10_5 (install instructions are here; build dependencies were installed by apt-get on chiara)
- ./configure --prefix=/ligo/apps/ubuntu12/nds2-client-0.10.5; make; make install
- In /ligo/apps/ubuntu12: ln -s nds2-client-0.10.5 nds2-client
nds2-client was apparently installed locally as a deb in the past, but the version in lscsoft seems broken currently (unknown symbols?). We should revisit this.
Built and installed pyepics (a cdsutils dependency)
- Download link to ~/src on chiara
- python setup.py build; python setup.py install --prefix=/ligo/apps/ubuntu12/pyepics-3.2.3
- In /ligo/apps/ubuntu12: ln -s pyepics-3.2.3 pyepics
pyepics was also installed as deb before; should revisit when Jamie gets back.
Added the gqrx ppa and installed gnuradio (dependency of the waterfall plotter)
Added a test in /ligo/apps/ligoapps-user-env.sh to load the new cdsutils only on Ubuntu 12.
The end result:
controls@chiara|~ > z
usage: cdsutils
Advanced LIGO Control Room Utilites
Available commands:
read read EPICS channel value
write write EPICS channel value
switch switch buttons in standard LIGO filter module
avg average one or more NDS channels for a specified amount of seconds
servo servos channel with a simple integrator (pole at zero)
trigservo servos channel with a simple integrator (pole at zero)
audio Play channel as audio stream.
dv Plot time series of channels from NDS.
water Live waterfall plotter for LIGO data
version print version info and exit
help this help
Add '-h' after individual commands for command help.
|
10187
|
Fri Jul 11 20:05:34 2014 |
Jenne | Update | LSC | QPD dark noise checkout |
The other day, I took spectra of the Yend transmission QPD dark noise for several different configurations of the transimpedance gain and the whitening gains.
I also calculated with Optickle the expected sensitivity vs. CARM offset for the sqrtInv CARM sensor.
I put these things together to get a plot of transmission QPD noise vs. CARM offset, for a chosen set of analog settings.
Measurement of Yend transmission QPD dark noise
Since EricQ had already checked out the whitening filters (see elog 9637 and elog 9642), I didn't check on them. I just left them (the analog whitening, and the digital antiwhitening filters) on.
First, I checked the noise vs. transimpedance gain. There are a few too many settings to put them all on one plot, so I have them sorted by the original transimpedance: 0.5 kOhms vs 5 kOhms. It's a little tricky to see, but all of the spectra that begin with the 5k transimpedance have a little extra noise around 10 Hz, although I don't know why. In the legend I have made note of what the settings were. x1 x1 is my representation of the "inactive" setting.

I then looked at the noise with different whitening gain slider settings. All but one of the traces are the 20 kOhm setting.

These .xml files are in /users/jenne/Arms/TransQPDnoise_July2014/
Calculation of inverse sqrt transmission sensitivity
I used Optickle to give me the power transmitted through the ETMs. I first find the transmission in the FPMI case. I use that to normalize the full PRFPMI transmission, so that the output units are the same as our C1:LSC-TR[x,y]_OUT units.
I take the square root of the transmitted power (sum of transmissions from each arm) at each CARM offset point, add 1e-3 as we do in the front end model to prevent divide-by-zero problems, and then take the inverse.
I find the slope by taking the difference in power between adjacent points, divided by the CARM offset difference between those points.
In this plot, I have taken the absolute value of the sensitivity, just for kicks. I also display an arbitrarily scaled version of the log of the transmitted power, so that we can see that the highest sensitivity is at half the maximum power.

Calculate the QPD dark noise in terms of meters
Finally, I put it all together, and find the dark noise of the QPD in terms of meters. Since the spectra were measured in units where the single-arm transmission is unity, the already match the units that I used to calculate the sqrtInv sensitivity.
I take the spectra of the QPD dark noise for the 20 kOhm case, and multiply it by the sensitivity calibration number at several different CARM offsets. As we expect, the noise is the best at half-max transmission, where the sensitivity is maximal.

|
10186
|
Fri Jul 11 17:49:12 2014 |
Nichin | Update | Electronics | New Prologix GPIB-Ethernet controller |
I have configured a NEW Prologix GPIB-Ethernet controller to use with HP8591E Spectrum analyzer that sits right next to the control room computers.
Static IP: 192.168.113.109
Mask: 255.255.255.0
Gateway: 192.168.113.2
I have no clue how to give it a name like "something.martian" and to update the martian host table (Somebody please help!!)
|
10185
|
Fri Jul 11 15:29:26 2014 |
Harry | Update | General | Telescope With Collimator |
I used a la mode to make a design for the coupling telescope with a 3.3um target waist, that included the collimator in the overall design. The plot is below, and code is attached.
The components are as follows:
label z (m) type parameters
----- ----- ---- ----------
lens1 0.7681 lens focalLength: 0.5000
lens2 0.8588 lens focalLength: 0.0350
collimator 0.8832 lens focalLength: 0.0020
The z coordinates are as measured from the beam waist of the NPRO (the figure on the far left of the plot).
Moving forward, this setup will be used to couple the NPRO (more specifically, the AUX lasers) light into the SM 980 fibers, as well as to help characterize the fibers themselves.
Ultimately, this will be a key component in the Frequency Offset Locking project that Akhil and I are working on, as it will transport the AUX light to the PSL, where the two beams will be beaten with each other to generate the input signal to the PID control loop, which will actuate the temperature servos of the AUX lasers. |
Attachment 1: telescope_2.zip
|
Attachment 2: telescopeWCollimator.png
|
|
10183
|
Fri Jul 11 11:51:03 2014 |
Nichin | Update | Electronics | PDFR: List of DC transimpedances |
The following values are going to be entered in the param_[PDname].yml file for each PD. I am elogging them for future reference.
I got the values from combing schematics and old Elog entries. Please let me know if you believe the values are different.
- AS55: 66.2 ohms
- REFL11 : 66.2 ohms
- REFL33 : 50.2 ohms
- REFL55: 50 ohms (Elog 4605)
- REFL165: 50.2 ohms
- POY11: 66.2 ohms
- POX11: 50.2 ohms
- REF (NF1611): 700 ohms
- POP22: ?? (This is currently a Thorlab BBPD )
|
10182
|
Fri Jul 11 09:41:43 2014 |
not Koji | Update | General | Coupling telescope design |
Quote: |
CFC-2X-C has a FIXED focal length of 2mm, but the collimator lens position is adjustable.
I'm not yet sure this affects your calculation or not as what you need is an approximate mode calculation;
once you couple the any amount of the beam into the fiber, you can actually measure it at the output of the fiber with a collimator attached.
|
I don't believe it had any effect, as all the calculations gave me the same target waist. |
10181
|
Fri Jul 11 08:11:56 2014 |
Steve | Update | CDS | ETMX needs help |
|
Attachment 1: ETMX.png
|
|
10180
|
Thu Jul 10 19:37:54 2014 |
Manasa | Update | General | PM980 fiber tested OK |
[Harry, Manasa]
This is the update from yesterday that Harry missed to elog.
We pulled out the first spool of the PM980 fiber yesterday and checked it using the illuminator at the SP table. Harry will be using this for all his tests and characterisation of the fiber. |
10179
|
Thu Jul 10 18:25:18 2014 |
Koji | Update | General | Coupling telescope design |
CFC-2X-C has a FIXED focal length of 2mm, but the collimator lens position is adjustable.
I'm not yet sure this affects your calculation or not as what you need is an approximate mode calculation;
once you couple the any amount of the beam into the fiber, you can actually measure it at the output of the fiber with a collimator attached. |
10178
|
Thu Jul 10 17:53:19 2014 |
Akhil | Configuration | Computer Scripts / Programs | EPICS installed on Raspberry Pi |
I finished the installation of EPICS base on Raspberry Pi (domenica: /opt/epics). I tested it by creating a test SoftIoc (controls@domenica~/freqCountIOC/simple.db) and was able to read from the channel on Chiara.
Now I am looking into how to call my C code that talks to Raspberry Pi through a .db script and write it into the assigned channel.
For installation I had to make these declarations in the environment (~/.bash_aliases):
export EPICS_EXT=${EPICS_ROOT}/extensions
export EPICS_EXT_BIN=${EPICS_EXT}/bin/${EPICS_HOST_ARCH}
export EPICS_EXT_LIB=${EPICS_EXT}/lib/${EPICS_HOST_ARCH}
if [ "" = "${LD_LIBRARY_PATH}" ]; then
export LD_LIBRARY_PATH=${EPICS_EXT_LIB}
else
export LD_LIBRARY_PATH=${LD_LIBRARY_PATH}:${EPICS_BASE_LIB}
fi
export PATH=${PATH}:${EPICS_EXT_BIN}
|
10177
|
Thu Jul 10 17:33:26 2014 |
jamie | Omnistructure | Computer Scripts / Programs | ubuntu12 software installed, gds 2.16.3.2 |
Rana wanted the latest GDS installed (for newest DTT), so I made an ubuntu 12 install directory into which I installed
- gds-2.16.3.2
- root_v5.34.03
I installed this stuff in
/ligo/apps/ubuntu12
which is the "official" location for stuff compiled specifically for ubuntu12.
Given that the workstations are diverging in OS (some ubuntu10, some ubuntu12), we're going to have to start supporting different software packages for the different versions, thus the new ubuntu12 directory. This will be a pain in the butt, and will certainly lead to different versions of things for different machines, different features, etc. We should really try to keep things at the same OS.
In any event, if you want to enable the GDS on an ubuntu 12 machine, source the ubuntu12 ligoapps-user-env.sh file:
controls@ottavia|~ > . /ligo/apps/ubuntu12/ligoapps-user-env.sh |
10176
|
Thu Jul 10 15:57:00 2014 |
Steve | Update | SUS | flow bench is turned off |
Flow bench effect on oplev error signal is here.
I turned off the south X-end flow bench. |
10175
|
Thu Jul 10 15:27:09 2014 |
Harry | Update | General | Coupling telescope design |
I designed this telescope to couple the 1064 NPRO into the PM980 fiber, using lenses from the Thorlabs LSB04-C kit.
The collimator is a CFC-2X-C, which has a variable focus length (2.0, 4.6, 7.5, and 11.0 mm) which gives corresponding angles of divergence of 0.298, 0.130, 0.79, and 0.054 degrees by the formula theta = (180*MFD) / (pi*f).
Then, using these values I calculated the spot size of a beam collimated by the CFC-2X-C, using f = w / tan(theta) where w is the spot size. This gave a value of 10.4 um.
I used this value (10.4 um) as a target waist for the telescope system, with the NPRO waist as a seed, at the origin.

It consists of two lenses, one located at Z = 77cm f = 50cm, and the second located at Z = 85.88 cm f = 2.54cm, which yields a waist of 13um at Z = 88.32cm, (which is where the collimator would go) for an overlap of .974.
Note that the telescope is so far "downrange" from the NPRO waist because it's a virtual waist, and the NPRO itself is located at about Z = 73cm.
Find attached the alm code used. |
Attachment 2: telescope.zip
|
10174
|
Thu Jul 10 02:15:36 2014 |
Jenne | Update | LSC | QPD dark noise checkout |
I have put beam dumps in front of both of the end transmission QPDs so that I could measure the dark noise. They are still there.
I checked that the Yend QPD sliders and switches were doing things as I expected. I couldn't do the Xend since c1auxex is still lost (elog 10165). I'll post plots and actual information on this checkout, as well as my calculation of what this dark noise means in terms of meters for CARM when we're using 1/sqrt(trans) signals tomorrow morning. |
10173
|
Thu Jul 10 02:09:20 2014 |
Jenne | Update | PSL | FSS Fast gain set |
I have put in a new nominal value for the FSS fast gain: 21.5 dB.
There is an oscillation peak in the MC error point spectra around 41.5 kHz if the FSS gain is set too high. I used the 4395 to have a look at the MC error point, and saw that if I set the FSS fast gain any lower than about 18 dB, the peak wasn't getting any smaller than -41 dBm. If I set the fast gain any higher than about 26 dB the peak wouldn't get any larger than about -34 dBm.
However, if I set the gain to 19.5dB, the PC RMS drive is consistently above 2 V, which isn't so good. If I crank the gain up to 27 dB or more, the PC RMS will stay below 0.9 V, which is great.
As a compromise, I have decided on 21.5 dB as the new FSS fast gain. This puts the oscillation peak at about -39.5 dBm, and the PC RMS around 1.6 V.
I changed the nominal gain by ezcawrite C1:PSL-STAT_FSS_NOM_F_GAIN 21.5 . This sets the nominal value so that the FSS screen's fast slider doesn't turn red at the new value. And, since the MC autolocker reads this epics channel and puts that into the gain during the mcup script, the MC autolocker now uses this new gain. For reference, it used to be set to 23.5 dB. |
10172
|
Thu Jul 10 01:02:13 2014 |
rana | Update | PSL | more PMC science |
Increased gain and SNR in PMC LO monitor circuit.
- R20: 499 -> 50k Ohms (increases gain by 100)
- Used Marconi to drive the LO input and readout C1:PSL-PMC_LODET
- Fit this function and loaded it into the psl.db file. The old Kalmus way used LOGE, but I wanted to use log10, so I did. The sensor is only useful in a narrow band. Since the signal is so low at low levels, I just fit to the highest 4 points because I was too lazy to do proper weighting. Do as I say, not as I do.
Plot with data and fit attached.
** N.B.: in order to update the calibration without rebooting, I used the following command: z write C1:PSL-PMC_LOCALC.CALC "2.235*LOG(B)+12.265". This allows us to update EPICS CALC records without rebooting the IOC. |
Attachment 1: PMCloCal.pdf
|
|
10171
|
Thu Jul 10 00:38:20 2014 |
manasa | Update | General | AOM and PSL Ringdown |
Quote: |
RXA: some more comments...
- The fact that the AOM can only modulate the power by a tiny bit means that it is very mis-aligned or that the driver is broken.
- You need to take into account the AOM step time in the calculation of the PMC step time. Its not a step response if the input step is not a step, but a exponential.
- I wouldn't trust that old John Miller entry for the PMC Finesse. As you can see from his elog, even he didn't trust it.
- As we were discussing before, making a little step is not the same as a full ringdown. cf. G000413 and T900007
|
I think we should revisit the AOM alignment because the last time it was aligned, PMC trans dropped from 0.84 to 0.15 (a little more than 80%) for 0-1V modulation input to the AOM driver [elog]. The drop in power right now is ~10-15% only.
I could not find any elogs of AOM alignment touchups after Oct 2012.
But can the ISS team throw some light on the status of AOM when they were installing the ISS servo before we decide on touching the AOM alignment? [elog] |
10170
|
Wed Jul 9 23:02:33 2014 |
Harry | Update | General | Beam Waists via Beamscan |
Today, I borrowed the beam profiler from Brian (another SURF) in order to double check my razor blade measurement figures, using the below setup:

Measurements are included in the a la mode code that is attached entitled beamfit.m. The beam fitting application yielded me waists (after the lens) of 35.44 um in the x plane, and 33.26 um in the y plane. These are both within 3 um of the measurements I found using the razor blade method. (I moved and resized the labels for the waists in the figure below for readability purposes.)

I then plugged these waists back into ALM, in addition to the lens specifications, to determine waist size and location of the NPRO, which turned out to be 543 um in the x located at Z = 1.160m, and 536 um in the y, located at 1.268m. These measurements are based upon zero at the waist after the lens, and the positive direction being back toward the NPRO.
 
The only systemic difference between these measurement and my original razor blade measurements was that I had taken the focal length of the lens as 75mm, which is advertised on the manufacturer's site. However, the more detailed specs revealed that the focal length was 85.8mm at 1064nm, which made a difference of about 400 um for the final waist determination. |
Attachment 4: beamScanWaist.zip
|
10169
|
Wed Jul 9 21:43:41 2014 |
Jenne | Update | Electronics | PMC servo card modifications in DCC |
[Rana, Jenne]
We have decided to keep better track (using new-fangled digital "computers") of our modifications to electronics boards.
The idea will be to create a new DCC document for every electronics board (when we pull a board and modify it, it should receive this treatment) that we have, and that document will become a history of the board's life. Version 1 will be a copy of the original drawing. Version 2 should be a modified version of that drawing with the current situation. All future versions should be modified from the most recent version, to reflect any changes. Notes for each updated version should include an elog reference to the work, so that we know why we did things, and have a place to find photos of the actual modifications. Elogs should also include a link to the DCC version. DCC titles should include the phrase "40m Revisions" for ease of searching.
Patient Zero for this new system will be the PMC servo card. The DCC number is D1400221. As of this moment, this just has the V1 original drawing with no modifications.
This has been included in the 40m's DCC document tree that Jamie started back in November 2012. |
10168
|
Wed Jul 9 21:05:31 2014 |
manasa | Update | General | AOM and PSL Ringdown |
After the fits, here are the numbers!
Component |
Measured |
Expected |
AOM |
85.1 ns |
200 ns (spec sheet) |
PMC |
164.6 ns |
Finesse/(2*pi*FSR) = 163.4 ns |
* We have a huge difference to the AOM switching time that was measured. The spec sheet mentions acoustic velocity in the material to be 4.2 mm/us and the well matched diameter in the AOM to be 1100 um. This would give a switching time ~ 200 ns. We could probably be having a much smaller beam size in the AOM for the measured switching time.
* The PMC parameters that I had been referring to from the wiki were actually wrong and which was the reason for the mismatch that I was finding. I modified the wiki according to the found references to the actual measurement here: PMC parameters The measured values now and then match pretty well.
* Since the AOM does not change the power of the output beam by very much, what we see is actually a step response. Also, we have a lot of noise in the data obtained at the PD.
RXA: some more comments...
- The fact that the AOM can only modulate the power by a tiny bit means that it is very mis-aligned or that the driver is broken.
- You need to take into account the AOM step time in the calculation of the PMC step time. Its not a step response if the input step is not a step, but a exponential.
- I wouldn't trust that old John Miller entry for the PMC Finesse. As you can see from his elog, even he didn't trust it.
- As we were discussing before, making a little step is not the same as a full ringdown. cf. G000413 and T900007
|
Attachment 1: data_code.zip
|
Attachment 2: PSL_ringFit.pdf
|
|
Attachment 3: AOMringFit.pdf
|
|
10167
|
Wed Jul 9 19:53:34 2014 |
rana | Update | PSL | PMC LO monitor trend (5 years) |

The first step is
The second uptick (In Nov 14, 2013) is when I removed a 3 dB attenuator from the LO line. Don't know why the decay accelerates after that. |
10166
|
Wed Jul 9 17:34:11 2014 |
Nichin | Update | Electronics | PDFR: Beam pointing adjustments and DC measurements |
[Nichin, Manasa]
AIM: Taking DC output readings with multimeter for each PD to create a database (required for transimpedance calculations), by taking off the table tops. Also, making sure each PD is illuminated properly.
What we did:
- In rack 1Y1: Diode laser controller was set to 150.0 mA at all times. This gave powers in the neighbourhood of 1mW at the end of fibers illuminating all PDs. The laser outputs light of 1064nm wavelength. The laser was switched off in the end.
- Checked the collimation of the fiber for each PD. In some cases they were not focused to give a sharp spot, so we had to unmount the fibers and fix it and mount them back. Manasa did it initially and I learnt how it was done properly. Eventually I got better and did it myself (under her supervision)
- Set the mount alignment for maximum illumination of the PD.
- Record the power falling on the laser and also the DC voltage output. Any light that did not come from my fiber was blocked when taking the readings and then unblocked. I also took care of offset voltage present when taking the DC readings.
Recorded measurements:
REFL11: Pinc = 0.91 mW VDC = 34.9 mV
REFL33: Pinc = 0.83 mW VDC = 33.2 mV
REFL55: Pinc = 1.08 mW VDC = 42.7 mV
REFL165: Pinc = 0.79 mW VDC = 115.3 mV
AS55: Pinc = 0.78 mW VDC = 31.3 mV
POX11: Pinc = 0.83 mW VDC = 34.7 mV
POP22**: Pinc = 1.08 mW VDC = 5.82 mV
POY11: Not illuminated; there was no optical fiber mount. Although, there was a fiber near it with a cap on the end. It also looks like there is no space to put in a new mount near the PD.
REF PD: Pinc = 1.19 mW VDC = 8.2 V (REF PD = New focus 1611)
**Note: The current POP 22 PD does not have 2 different outputs for DC and RF signals. I unplugged the RF cable from the output, took readings with the multimeter and then plugged back the RF cable.
Further work:
I will calculate the responsivity for each PD and compare it to the expected values. |
10165
|
Wed Jul 9 17:08:29 2014 |
Jenne | Update | CDS | c1auxex "Unknown Host" |
c1auxex has forgotten who it is. Slow sliders for the QPD head were not responding, so I did a soft reboot from telnet. The machine didn't come back, so I plugged the RJ45-DB9 cable into the machine and looked at it through a minicom session. When I key the crate, it gives me an error that it can't load a file, with the error code 0x320001. Looking that up on a List of VxWorks error codes, I see that it is: S_hostLib_UNKNOWN_HOST (3276801 or 0x320001)
I'm not sure how this happened. I unplugged and replugged in the ethernet cable on the computer, but that didn't help. Rana is going in to wiggle the other end of the ethernet cable, in case that's the problem. EDIT: Replacing the ethernet cable did not help.
Former elogs that are useful: 10025, 10015
EDIT: The actual error message is:
boot device : ei
processor number : 0
host name : chiara
file name : /cvs/cds/vw/mv162-262-16M/vxWorks
inet on ethernet (e) : 192.168.113.59:ffffff00
host inet (h) : 192.168.113.104
user (u) : controls
flags (f) : 0x0
target name (tn) : c1auxex
startup script (s) : /cvs/cds/caltech/target/c1auxex/startup.cmd
Attaching network interface ei0... done.
Attaching network interface lo0... done.
Loading...
Error loading file: errno = 0x320001.
Can't load boot file!!
|
10164
|
Wed Jul 9 16:33:05 2014 |
manasa | Update | PSL | PMC ringdown setup |
Quote: |
Quote: |
I moved stuff on the PSL table to accommodate the PMC ringdown setup.
I used the beam that leaks from the steering mirror at the PMC transmission that was dumped to a razor blade dump. I installed a Y1 to steer the beam to the ringdown PD. Power in the beam 75mW.
|
I am guessing that 75 mW will burn / destroy any Thorlabs PD. I hope that mW is supposed to be uW.
|
It was ~7.5mW and measured ~2V at the PD output (given its range 0-5V ) on the oscilloscope . So PD is safe ! |
10163
|
Wed Jul 9 12:01:43 2014 |
Akhil | Configuration | Electronics | Setup Plan for placing the Frequency Counter inside the lab |
Today, me and Manasa went inside the lab to figure out a place for the place for the FC. The whole setup will be placed in a chassis box . The chassis in figure(setupforFC.pdf) will be placed in the highlighted(red) box in the figure(setup.png). All the cables will be routed to the computers from behind the box and the RF cables from the beat box will be routed from the front end of the box. The two raspberry Pi boxes will be placed inside the box and the Frequency counters will be mounted as shown so that the frequency count can be seen from outside. |
Attachment 1: setup.png
|
|
Attachment 2: SetupFC.png
|
|
10162
|
Wed Jul 9 11:41:12 2014 |
Koushik | Update | PSL | PMC local oscillator is going wonky |
Quote: |
Koushik replaced an ERA-5 in the PC path. We put the module back to the rack and found no change.
The epics LO level monitor monitor is still fluctuating from 6~11dBm. We need more thorough investigation
by tracing the signals everywhere on the board.
Despite the poor situation of the modulation, PMC was locking (~9PM). Rana suspect that the PMC demodulation
phase was not correctly adjusted long time.
Koushik has the measured power levels and the photos of the board. I'll ask him to report on them.
|
Updates from Koushik:
The power levels measured (before and after relacement of ERA-5) are as follows:
LO to Servo : Vout = 2.3 Vpp / Pout = 11.21 dBm at f = 35.5 MHz
RF to PC : Vout = 354 mVpp / Pout = -5.1 dBm at f= 35.5 MHz
The measurements were done using an oscilloscope with 50 ohms load impedance. Unfortunately the photos are not available from the camera. |
10161
|
Wed Jul 9 08:50:30 2014 |
Akhil | Update | General | Weekly Update |
Last Week's Work:
- Worked on the setup to mitigate timing issues arising due to the non-synchronization of clocks of the Frequency counter and Raspberry Pi .
- Took measurements with the mentioned setup and generated PSD plots of the FC.
- Completed the setup for phase measurements by using an external ADC.
Work Plan for this Week:
- Complete the installation of the Mini Circuits Frequency Counter on the EPICS. This involves installation of EPICS base on Raspberry Pi, creating an IOC server on the R Pi and writing the data from the FC into a specific IOC channel.
- Complete phase measurements and obtain the delay in the FC thus completing the characterization of the FC.
- Install the FC at a suitable place inside the 40m so that the beat note system can be remotely managed from any of the Control computers thus effectively replacing the spectrum analyzer(This will be done with proper supervision once the recently ordered FC is shipped).
Inside the 40m :
- I will be going inside the lab today around 9 am with Manasa to make a plan about where the FC must be placed and the routing of the RF cables and the cables which run into computers from the FC.
- Once the channel is created and tested, we will install the FC inside the lab possibly by the end of this week or by next week.
|
10160
|
Wed Jul 9 00:59:09 2014 |
rana | Update | PSL | PMC local oscillator is going wonky |
Quote: |
Koushik and Koji try to fix the PMC oscillator issue. So we remove the module from the rack.
This means we don't have the PMC transmission during the work.
|
After the ERA-5 was replaced (see Koushik elog) we relocked the PMC.
The new LO level going into the PMC servo card is +11.5 dBm. The LO mon on the PMC card reads 9 dBm and seems so flat I now suspect the monitor circuit.
I also measured the RF drive to the EOM as a function of C1:PSL-PMC_RFADJ on the Phase Shifter screen.
The phase shifter slider gives ~75 deg/V in phase shift of the RF out to the EOM. I tried to optimize the loop gain quickly using the fluctuations in the reflected power. The loop oscillates at high frequency with the slider at 21 dB and also at 9 dB. So I set the gain at +14 dB. Needs to be optimized correctly in the daytime.
|
Attachment 1: PMC_RFslider.pdf
|
|
10159
|
Wed Jul 9 00:47:22 2014 |
Koji | Update | PSL | PMC local oscillator is going wonky |
Koushik replaced an ERA-5 in the PC path. We put the module back to the rack and found no change.
The epics LO level monitor monitor is still fluctuating from 6~11dBm. We need more thorough investigation
by tracing the signals everywhere on the board.
Despite the poor situation of the modulation, PMC was locking (~9PM). Rana suspect that the PMC demodulation
phase was not correctly adjusted long time.
Koushik has the measured power levels and the photos of the board. I'll ask him to report on them. |
10158
|
Tue Jul 8 23:59:49 2014 |
Harry | Update | General | Weekly Plan (7.8.14) |
Last Week:
-I continued to struggle with the razorblade beam analysis, though after a sixth round of measurements, and a lot of fiddling around with fit parameters in matlab, there seems to be a light at the end of the tunnel.
Next Week:
-I plan to check my work with the beamscan tomorrow (wednesday) morning
-Further characterize the light from the fibers, and set up the collimator
-Design and hopefully construct the telescope that will focus the beam into the collimator
Materials:
- Razorblade setup or beamscan (preferably beamscan)
- Fiber Illuminator
- Collimator (soon to be ordered)
- Lenses for telescope (TBD)
|
10157
|
Tue Jul 8 22:53:02 2014 |
Jenne, Rana | Update | Electronics | Transmon QPD / whitening |
We need to work farther on checking out the end transmission QPD electronics situation.
In bullet-point form, we need to:
* Ensure that the Weiss QPD head modifications have been made on these diodes. (cf. Rai W's LLO elogs on QPDs)
* Ensure that the QPD biases are somewhere in the range of 10-15V, not the old 100V. (Because we only need HV to make the capacitance low for RF use. Low voltage means less power dissipation in the head)
* Ensure the Rana/Rob modifications have been propagated to the whitening boards, so that we have full dynamic range. (Steve is looking for the marked up paper schematics)
* Replace signal path resistors with low noise metal film resistors.
* Check QPDs / whitening boards for oscillation (with a scope probe), ensure that we chose an appropriate analog gain.
In thinking about the transimpedances that we want, we thought about the current that we expect. We should get about 100 mW of light transmitted through the ETMs when we have full IFO lock. There is a 50/50 BS to split the light between the QPD and the Thorlabs transmission diode, so we have about 50 mW incident on the QPDs, which is about 13 mW per quadrant. With a sensitivity of about 0.15 Amps/Watt for silicon, this means that we're expecting to see about 2 mA of current per quadrant once we have the IFO fully resonant. We want this to correspond to about 5V, which means we want a transimpedance gain of around 2.5 kOhm.
For the things that need checking, each quadrant has:
Photodiode ------ Gain Switch 1 ----- Gain Switch 2 ------ Gain Switch 3 ------ Variable Gain Amplifier ------- Whitening stage 1 (z @ 4 Hz, p @ 40 Hz) ------- Whitening stage 2 (z @ 4 Hz, p @ 40 Hz)
We want to check on the status of each of these switches, and whether they actually do what they say on the QPD Head screens. Q has checked out and fixed the bit outputs for the whitening stages, but the rest still needs to be checked out. Also note that the Switch 1, Switch 2 and Switch 3 are common to all 4 quadrants (i.e. enabling switch 1 on one quadrant enables it on all quadrants), but the variable gains and the whitening stages are individual for each quadrant. |
10156
|
Tue Jul 8 18:20:15 2014 |
jamie | Omnistructure | Electronics | Jamie 1811 power supply fixed! |
Placed in PD cabinet in Y arm, next to the OTHER Jamie-made 1811 power supply from 1998. |
10155
|
Tue Jul 8 17:59:12 2014 |
jamie | Omnistructure | Electronics | Jamie 1811 power supply fixed! |
I finally made good on the LIFE TIME WARRANTY on the ancient, Jamie-made 1811 power supply with the faulty switch:

Back to fully working form. Hopefully I'll still be around the next time it breaks in 16 years. |
10154
|
Tue Jul 8 16:45:15 2014 |
Nichin | Update | General | Weekly plan |
My plan for next week is...
1) 1) Taking DC output readings with multimeter for each PD to create a database for all the PDs. Requires taking off the table tops for each PD. Also, making sure each PD is illuminated properly.
2 - 3 Hours inside the lab
Requires presence of expert
Occupies all the PDs , RF switch and the Network analyzer.
2) 2) Integrate the switch selection script with the Network analyzer script to complete the automation part of the project. (If time permits, build a simple GUI for easy operation)
Occupies the control room computer, RF switch and the Network analyzer
3) 3) Create a database of canonical plots for each PD to compare with the current plot and maybe even plot the difference between the current plot and canonical plot.
Occupies the control room computer, PDs , RF switch and the Network analyzer.
4) 4) Fit the transfer function or transimpedance using vector fitting. (vectfit4.m)
5) 5) Update 40m-Wiki
6) 6) Progress Report to be submitted to SFP. |
10153
|
Tue Jul 8 15:28:32 2014 |
Koji | Update | PSL | PMC local oscillator is going wonky |
Koushik and Koji try to fix the PMC oscillator issue. So we remove the module from the rack.
This means we don't have the PMC transmission during the work. |
10152
|
Tue Jul 8 15:07:24 2014 |
Nichin | HowTo | Electronics | RF Multiplexer in rack 1Y1 |
The RF multiplexer is configured as shown in the figure. It is now effectively a 15x1 RF mux.

To select a required channel:
Run the script as shown below
/opt/rtcds/caltech/c1/scripts/general/rfMux.py
>python rfMux.py ch11
For channel 10 to 16, you can just enter the required channel number and it is routed to the output.
For channel 1 to 8, you only need to input the required channel number as above. No need to run the code again to select ch9 after selecting ch1-8
How the NI-8100 controller works:
Whenever any channel of one switch is selected, the output of the other switch is set to its ch0 (ch1 and ch9 in the figure).
So selecting ch1-8 will automatically select ch9 as output for the other switch. IF you send a command to select ch9 afterwards, the first switch will be automatically set to ch1 and not stay on what you had selected before. |
10151
|
Tue Jul 8 09:02:02 2014 |
Akhil | Update | Electronics | PSD Plots for different sampling times of the Frequency Counter |
Although there were few timing issues with the FC and the Raspberry Pi at the lowest sampling time of the FC (0.1s) even after adding an external trigger circuit, it turned out that most of these issues are not prevalent at higher sampling times(>0.5 s)(narrow peaks of PSD seen for higher sampling times). Rana suggested me to look at the PSD plots at different sampling times of the FC so that we can decide which would be the optimal sampling time to work with the FC before replacing the spectrum analyzer. I took the measurements with the setup discussed in my previous elog(http://nodus.ligo.caltech.edu:8080/40m/10129) . However, the noise of the R Pi- FC interface should be taken care of (I will discuss it with my mentors).
Attached are the plots at 100 MHz carrier frequency at different sampling times of the FC( 0.1s, 0.2s, 0.3s, 0.5s, 1s) (pdfs and code attached in a zip file)
RXA: Put all the plots in a single PDF file and use the same axis limits for all plots so that its easy to compare. (Attached in PSD.pdf)
|
Attachment 1: 0.1s.png
|
|
Attachment 2: 0.2s.png
|
|
Attachment 3: 0.3s.png
|
|
Attachment 4: 0.5s.png
|
|
Attachment 5: 1s.png
|
|
Attachment 6: Pdf.zip
|
Attachment 7: PSD.pdf
|
|
10150
|
Tue Jul 8 01:55:21 2014 |
Jenne | Update | SUS | ETMX glitching |
[Jenne, Rana]
A few times this evening, I had been having trouble locking CARM and DARM with ALS, and holding it for very long. When it started happening again, I switched over to locking the individual arms with ALS. Yarm seems to be totally fine, but Xarm has something funny going on.
Rana and I have narrowed it down to being a problem with ETMX. We were watching ETMX's oplev and local damping error signals, and would see occasional glitch events. This happened when oplev + local damping were both on, both off, and when only local damping was on. We believe that this points to something weird with the coil driver and actuator chain.
We tried to watch for a while to see if it was a step event (something switching on and off periodically), or an impulse event (some transient oscillation in an opamp perhaps), but the problem went away again. We have come to no conclusions other than we have a problem that needs watching.
During our investigations, to more softly turn off the damping, Rana set the local damping gains, as well as the oplev gains to zero using a ramp time. We don't recall the precise numbers, and conlog doesn't have the gains recorded, so we made an educated guess. The local damping seems fine, but the oplev damping should be re-confirmed. Steve, can you please show Harry how, and have him help you measure the ETMX pitch and yaw oplev loops, and set the gains so that they match up to the references, and then post the measured bode plots when you're done? |
10149
|
Mon Jul 7 23:19:55 2014 |
rana | Update | PSL | PMC ringdown setup |
Quote: |
I moved stuff on the PSL table to accommodate the PMC ringdown setup.
I used the beam that leaks from the steering mirror at the PMC transmission that was dumped to a razor blade dump. I installed a Y1 to steer the beam to the ringdown PD. Power in the beam 75mW.
|
I am guessing that 75 mW will burn / destroy any Thorlabs PD. I hope that mW is supposed to be uW. |
10148
|
Mon Jul 7 22:18:26 2014 |
Koji | Update | PSL | PMC local oscillator is going wonky |
It seems that there is no better chip in MiniCircuits line-up with the same form factor.
ERA-5 is the most powerful one in the ERA (or MAR) series.
If the output is ~0dBm we have MAR-6SM in stock. But I suspect that ERA-5 was driven at the power level close to its saturation (~18dBm).
If we allow different form factors, we have GVA-** or GALI-** in the market and also in the blue tower, in order to gain more performance margin.
If it is difficult to apply them, I would rather use another ERA-5 with enhanced heat radiation.
I'm sure that Downs has EAR-5 replacement. |
10147
|
Mon Jul 7 22:07:29 2014 |
rana | Update | Electronics | RF cables rerouted |
This Red Ladder movement also probably included some bumping of the MC2 stack (be more careful when working in the lab). Around 5 PM today, the MC lost lock.
When I came in later, I found that the MC was flashing the TEM17,9 mode and had been misaligned like this for ~1 hour. I looked through the MC SUS sensor trends and moved MC2 back to its old position to get the locking back. I disabled the WFS, tweaked the TEM00 mode alignment using only MC2 sliders, and then re-enabled WFS. Its been OK for the past 5 hours. |
Attachment 1: Untitled.pdf
|
|